OpenStudy (jolo_yolo):

A function is shown below: f(x) = (0.93)^x What does the function represent? Exponential growth of 7% Exponential decay of 7% Exponential growth of 93% <--- correct? Exponential decay of 93%

OpenStudy (jolo_yolo):

So so what's the answer then, A? it's exponential, which means increasing no decaying right? @uhhhhhh

OpenStudy (perl):

exponential can decrease

OpenStudy (anonymous):

Growth or decay depends on whether what's inside the parenthesis is greater or less than one. Since .93 is less than one, it has to be decay.

OpenStudy (anonymous):

Exponential decay of 7% since 0.93 - 93%

OpenStudy (perl):

in general the expnential function is y = A ( 1 + r ) ^x if r is negative then it is decay. if r is positive then it is growth.

OpenStudy (perl):

or y = A ( 1 + r/100 ) ^ x

OpenStudy (anonymous):

The function is measuring on the decay of f(x) = < 100%. That's how I interpret it.

OpenStudy (jolo_yolo):

oh okay i see, final answer B, correct?

OpenStudy (anonymous):

Just to clarify my thought process: Since if you graph f(x) = 1^x and f(x) = 0.93^x side by side, it presents the exponential decay of 7%.

OpenStudy (anonymous):

but then again it might be growth, I could be wrong, cause the curve is upwards and not downwards

OpenStudy (jolo_yolo):

ightt, imma just go with decay of 7%

OpenStudy (anonymous):

lemme refer the question first pls

OpenStudy (jolo_yolo):

okay

OpenStudy (anonymous):

@saifoo.khan

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

I think you should stick with C

OpenStudy (jolo_yolo):

Okee, well thanks for ur help